#4 Derek plans to buy a $30,976.00 car. The dealership offers zero percent financing for 53.00...

#4
Derek plans to buy a $30,976.00 car. The dealership offers zero percent financing for 53.00 months with the first payment due at signing (today). Derek would be willing to pay for the car in full today if the dealership offers him $____ cash back. He can borrow money from his bank at an interest rate of 5.04%.


Submit
Answer format: Currency: Round to: 2 decimal places.
Answer #1

Basic Value of the Car = $30,976.

No of months for which  zero percent financing is offered = 53 months

Since, the first payment is due at signing (today) , we will calculate the PVIFA for 52 months to arrive at the present value of the car

Monthly payment required to be made = 30,976/53 = $584.4528

Interest rate per month = 5.04%/12 = 0.42%

Present Value of the Car at zero cost financing for 53 months with the first payment due at signing (today) = Monthly payment required to be made*PVIFA(0.42%,52)

= 584.45288*46.6258

= $27,250.58

.

Derek would be willing to pay for the car in full today if the dealership offers him the following cash back:

Basic Value of the Car - Present Value of the Car at zero cost financing for 53 months with the first payment due at signing (today)

= 30,976 - 27,250.58

= $3,725.42